3 1/4y - 7 1/2 = 2 2/3
7/x - 2 = 4/x + 3
Answer by Fruglemiester1234(13) 
 I assume you are just trying to find X and Y in these equations.
1) For this problem I would first of all get the mixed fractions into single fractions(not sure their specific name) Like:
13/4y - 15/2 = 8/3 note: you can do this by times the outside # by the denominator then adding the numerator....3 1/4 ... 3*4 +1 =13 ...13/4
Then I would get them all to have the same denominator(bottom #)...
39/12y - 90/12 = 32/12 note: the least common denominator for 4,3,2 is 12. Times both the 13 and the 4 by 3, you get 39 for the top and 12 for the bottom, do this for all three.
Add 90/12 to both sides. You get:
39/12y = 32/12 + 90/12
You can add 32 and 90, because they both have the same bottom #. you get:
39/12y = 122/12
Divide both sides by 39/12. Instead of dividing on the other side you can times by the recipricol, as follows:
y = 122/12 / 39/12 ... y = 122/12 * 12/39... y = 122*12/39*12
12's end up cancelling, because theres one on the top and one on the bottom (like 12/12 = 1). After this you would try to reduce the fraction of 122/39, but in this case it is as simple as it can get.
y = 122/39
or if you want decimal
y = 3.13
2) 7/x - 2 = 4/x +3
First I would try to get the x's together on one side. We can start to do this by adding 2 to both sides. we end up with:
7/x = 4/x + 5
Then I would subtract (4/x) from both sides.
7/x - 4/x = 5
Sense both fractions have the same denominator we can add(in this case subtract) the numerators and write them as one fractions, like this:
(7-4)/x = 5 ...7 subtract 4 gets us to..
3/x =5
Then we times both sides by x, getting the x out from beneath the 3.We get:
3 = 5x
From here its pretty basic, we divide by 5, getting us:
3/5 = x
X = 3/5
or in decimal value
X = .6
| |
|